OpenStudy (anonymous):

what is the solution to the system below 2x-y=2 2x-y=-2 infinite solutions on the line (2, -2) no solution (2, 2)

OpenStudy (anonymous):

2x - y = 2 --->(-1) 2x - y = -2 ----------------- -2x + y = -2 (result of multiplying by -1) 2x - y = -2 ----------------add 0 = -4 (incorrect) NO SOLUTIONS

OpenStudy (anonymous):

They are parallel lines (they have the same slope) therefore they will never meet. So no sol'n.